# The 2-systole on compact Kähler surfaces with positive scalar curvature

### v1

@@ 4.1. Geometry of ruled surfaces and notation
Let $X_{0}\to B$ be a ruled surface over a smooth complex curve $B$ of genus $g$. Denote by $C_{0}$ a minimal section and by $F$ the fibre class in $X_{0}$, satisfying
$F^{2}=0,\quad C_{0}\cdot F=1,\quad C_{0}^{2}=-e,\quad e\in\mathbb{Z}_{\geq 0}.$
The first Chern class and a Kähler class of $X_{0}$ are given by
$c_{1}(X_{0})=2C_{0}+(2-2g+e)F,\qquad[\omega]=aC_{0}+bF,\quad a>0,\ b>ae.$
A direct computation shows that
@@
Let $X_{0}$ be a ruled surface over a smooth complex curve $B$ of genus $g$, with invariant $e\geq 0$. Then
@@
Moreover, when $g=0$ the supremum is attained precisely on classes proportional to $C_{0}+(e+1)F$, and when $g\geq 1$ the supremum is attained if and only if $g=1$ and $b/a\geq e+1$.
@@
It is well known that the Mori cone $\overline{\mathrm{NE}}(X_{0})$ is generated by the numerical classes of the fibre and a minimal section, in our notation, this is $F$ and $C_{0}$. Every other effective curve is numerically equivalent to $C_{0}+nF$ with $n\geq 0$, and has intersection at least as large with $\omega]$ as one of these generators. Hence
@@
The quantity $\mathcal{J}_{X_{0}}$ is homogeneous of degree 0 in $\omega]$, so it depends only on the ratio $t:=b/a$. The Kähler condition $b>ae$ becomes $t>e$. Writing
$\mathcal{J}_{X_{0}}([\omega])=4\pi\,\phi_{e,g}(t),\quad\text{with}\quad\phi_{e,g}(t):=\min\{1,\ t-e\}\cdot\frac{2t+2-2g-e}{2t-e},\qquad t>e,$
it suffices to compute $\sup_{t>e}\phi_{e,g}(t)$. We distinguish two cases according to $g$.
Case 1: $g\geq 1$. For $g\geq 1$ we have $2-2g\leq 0$, hence
$2t+2-2g-e\;\leq\;2t-e,$
and therefore
$\frac{2t+2-2g-e}{2t-e}\leq 1,\qquad t>e.$
It follows that
$\phi_{e,g}(t)=\min\{1,\ t-e\}\cdot\frac{2t+2-2g-e}{2t-e}\leq\min\{1,\ t-e\}\leq 1,$
so $\phi_{e,g}(t)\leq 1$ for all $t>e$.
If $g=1$, then 22ge=e2-2g-e=-e and
$\phi_{e,1}(t)=\min\{1,\ t-e\}\cdot\frac{2t-e}{2t-e}=\min\{1,\ t-e\}.$
Hence $\sup_{t>e}\phi_{e,1}(t)=1$, and the supremum is attained precisely when $t-e\geq 1$, i.e. $b/a\geq e+1$.
If $g>1$, then 22g<02-2g<0 and the inequality 2t+22ge<2te2t+2-2g-e<2t-e is strict, so
$\phi_{e,g}(t)<\min\{1,\ t-e\}\leq 1,\qquad t>e.$
On the other hand,
$\lim_{t\to+\infty}\phi_{e,g}(t)=\lim_{t\to+\infty}\frac{2t+2-2g-e}{2t-e}=1,$
so $\sup_{t>e}\phi_{e,g}(t)=1$, but it is not attained for any finite $t$. Consequently, in all cases $g\geq 1$ we obtain
$\sup_{t>e}\phi_{e,g}(t)=1,\quad\text{and}\quad\sup_{\mathcal{K}(X_{0})}\mathcal{J}_{X_{0}}([\omega])=4\pi.$
Case 2: $g=0$. Now 22ge=2e2-2g-e=2-e, and for $t>e$ we split
For $t\geq e+1$, a straightforward derivative computation gives
$\phi_{e,0}^{\prime}(t)=\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{2t+2-e}{2t-e}\right)=-\frac{4}{(2t-e)^{2}}<0,$
so on $e+1,+\infty)$ the function $\phi_{e,0}$ is strictly decreasing, and
$\max_{t\geq e+1}\phi_{e,0}(t)=\phi_{e,0}(e+1)=\frac{2(e+1)+2-e}{2(e+1)-e}=\frac{e+4}{e+2}.$
For $e<t\leq e+1$, a direct computation shows that $\phi_{e,0}^{\prime}(t)>0$ for all $t>e$, so $\Phi_{e,0}$ is strictly increasing on $(e,e+1]$ and
$\max_{e<t\leq e+1}\phi_{e,0}(t)=\phi_{e,0}(e+1)=\frac{e+4}{e+2}.$
Combining the two ranges, we see that $\phi_{e,0}$ attains its global maximum on $(e,+\infty)$ at $t=e+1$, with value
$\sup_{t>e}\phi_{e,0}(t)=\phi_{e,0}(e+1)=\frac{e+4}{e+2}.$
This corresponds exactly to classes proportional to $C_{0}+(e+1)F$. Thus, in the case $g=0$,
$\sup_{\mathcal{K}(X_{0})}\mathcal{J}_{X_{0}}([\omega])=4\pi\,\frac{e+4}{e+2},$
and the supremum is attained precisely on Kähler classes with $b/a=e+1$. ∎
@@
It is well known that a ruled surface over the projective line is a Hirzebruch surface $\mathbb{F}_{e}$ with invariant $e\in\mathbb{Z}_{\geq 0}$. When $e=1$, $\mathbb{F}_{1}$ is not minimal and is isomorphic to the blowup $\operatorname{Bl}_{1}\mathbb{P}^{2}$ of $\mathbb{P}^{2}$ at one point. As expected, our computations of $\mathcal{J}_{\operatorname{Bl}_{1}\mathbb{P}^{2}}$ in Example 3.5 and of $\mathcal{J}_{\mathbb{F}_{1}}$ in Proposition 4.2 yield the same value.
For later use, we also fix notation for blowups of ruled surfaces. We do assume that all blowup points are in very general positions, that is, each point lies on a distinct fiber of the ruling and that no point lies on the negative section $C_{0}$. Indeed, if some points collide on the same fiber or lie on $C_{0}$, additional effective curves appear whose strict transforms have smaller intersection with any fixed Kähler class, so they can only decrease $\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$ and hence $\mathcal{J}_{X_{k}}([\omega])$. Thus this configuration is the worst case for the supremum.
Let $X_{k}=\text{Bl}_{k}X_{0}$ be the blowup of $X_{0}$ at $k$ points in very general position, and denote by $E_{1},\dots,E_{k}$ the exceptional divisors. Then
@@
$C_{0}^{2}=-e,\quad F^{2}=0,\quad C_{0}\cdot F=1,\quad E_{i}^{2}=-1,\quad E_{i}\cdot C_{0}=E_{i}\cdot F=E_{i}\cdot E_{j}=0\ (i\neq j).$
A Kähler class on $X_{k}$ can be written as
$\omega]=aC_{0}+bF-\sum_{i=1}^{k}t_{i}E_{i},\qquad a>0,\ b>ae,\ t_{i}>0,$
and the first Chern class is
@@ A Kähler class on $X_{k}$ can be written as
$\omega]^{2}=2ab-ea^{2}-\sum_{i=1}^{k}t_{i}^{2},\qquad c_{1}(X_{k})\cdot[\omega]=2b+(2-2g-e)a-\sum_{i=1}^{k}t_{i}.$
These expressions will be used repeatedly in the subsequent subsections.
@@ 4.2. Non-rational ruled surfaces
Let $X_{0}$ be a minimal ruled surface over a smooth curve $B$ of genus $g\geq 1$, with invariant $e\geq 0$. For an integer $k\geq 1$, let $X_{k}=\text{Bl}_{k}(X_{0})$ be the blow-up at $k$ points in very general position, with exceptional divisors $E_{1},\dots,E_{k}$. Then for every $k\geq 1$ one has
@@
Since $\mathcal{J}_{X_{k}}([\omega])$ is homogeneous of degree 0 in $\omega]$, we may therefore normalize $a=1$ and write
$\omega]=C_{0}+bF-\sum_{i=1}^{k}t_{i}E_{i},\qquad b>e,\ t_{i}>0,$
with the Kähler condition implying $2b-e-\sum t_{i}^{2}>0$. As in the minimal case, $\mathcal{J}_{X_{k}}([\omega])$ can be written as
@@
On the other hand, using $g\geq 1$ and $0<u_{i}<1$ we have
@@
It remains to show that the bound $2\pi$ is sharp. Consider the family of Kähler classes
$\omega_{b}]=C_{0}+bF-\sum_{i=1}^{k}t_{i}E_{i},\qquad b>e.$
In fact, any effective curve on $X_{k}$ has the form $C=dC_{0}+nF-\sum m_{i}E_{i}$ for $d,n,m_{i}\in\mathbb{Z}$, with the intersection number $\omega\cdot C=(b-e)d+n-m_{i}t_{i}$. We take $b$ large enough, so that the curves $E_{i}$ or $F-E_{i}$ realize the holomorphic $2$-systole, i.e.
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{b}])=\min_{1\leq i\leq k}\{t_{i},1-t_{i}\}.$
Observe that, for all $b>0$ large enough, the classes $\omega_{b}]$ remain in the Kähler cone, and
$\hat{S}([\omega_{b}])=4\pi\frac{2b+2-2g-e-\sum t_{i}}{2b-e-\sum t_{i}^{2}}\nearrow 4\pi,\qquad\text{as}\quad b\nearrow\infty.$
Therefore, by setting $t_{i}=1/2$, we have
$2\pi\frac{2b+2-2g-e-k/2}{2b-e-k/4}=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{b}])\cdot\hat{S}([\omega_{b}])=\mathcal{J}_{X_{k}}([\omega_{b}])\leq\sup_{\mathcal{K}^{+}(X_{k})}\mathcal{J}_{X_{k}}\leq 2\pi.$
Taking $b\nearrow\infty$, we arrive at our conclusion. ∎
@@
Moreover, equality holds if and only if $B$ is an elliptic curve, $X$ is minimal and isometrically covered by $\mathbb{P}^{1}\times\mathbb{C}$, and $\omega$ has constant scalar curvature so that $\omega\cdot F\leq\omega\cdot C_{0}$, in which case the $\operatorname{sys}_{2}(\omega)$ is achieved by the $\mathbb{P}^{1}$-fibre.
@@
By Proposition 4.2 and Lemma 4.4, we obtain
$\operatorname{sys}_{2}(\omega)\cdot\min_{X}S(\omega)\leq\sup_{\mathcal{K}^{+}(X)}\mathcal{J}_{X}([\omega])=4\pi,$
with equality if and only if $B$ is an elliptic curve, $\omega$ has constant scalar curvature, and the holomorphic $2$-systole is realised by the $\mathbb{P}^{1}$-fibre. It therefore suffices to work in this equality case and prove that $X$ is isometrically covered by $\mathbb{P}^{1}\times\mathbb{C}$.
Since $X\xrightarrow{\pi}B$ is a minimal ruled surface fibred over an elliptic curve, there exists a rank $2$ holomorphic vector bundle $\mathcal{E}$ on $B$ such that $X=\mathbb{P}(\mathcal{E})$ and $\pi$ is the bundle projection (see [Har77, Chap. V]). By the classification of cscK metrics on ruled surfaces due to Apostolov and Tønnesen-Friedman (see [ATF06, Thm. 2], the existence of a cscK metric on $X=\mathbb{P}(\mathcal{E})$ is equivalent to the slope-polystability of the underlying rank $2$ bundle $\mathcal{E}$ over the base curve $B$. Hence our cscK metric $\omega$ forces $\mathcal{E}$ to be polystable, and we may write $\mathcal{E}=\bigoplus_{i}\mathcal{E}_{i}$ as a direct sum of stable bundles of the same slope. By Fujiki’s equivalences [Fuj92, Lemma 2], this implies that $X=\mathbb{P}(\mathcal{E})$ is quasi-stable. Equivalently, there exists a projective unitary representation
@@
where $\widetilde{B}\cong\mathbb{C}$ is the universal cover and $\pi_{1}(B)$ acts diagonally. Fix a biholomorphism $\Phi:X\stackrel{{\scriptstyle\sim}}{{\to}}X_{\rho}$ and define
@@
Then $\omega_{\rho}$ is a cscK Kähler metric on $X_{\rho}$, and by construction $\Phi$ becomes an isometry between $(X,\omega)$ and $(X_{\rho},\omega_{\rho})$.
Since $g(B)=1$, Fujiki’s rigidity result [Fuj92, Lemma 10] shows that any cscK Kähler metric on $X_{\rho}$ is a generalized Kähler–Einstein metric in the sense of [Fuj92, §3]. This yields $X_{\rho}$ is isometrically covered by $\mathbb{P}^{1}\times\mathbb{C}$ endowed with the product Kähler metric $\omega_{\text{flat}}\oplus\omega_{\text{FS}}$, so is $X$. ∎

### v2

@@ 4.1. Geometry of ruled surfaces and notation
Let $X_{0}\to B$ be a minimal ruled surface over a smooth complex curve $B$ of genus $g$. Let $F$ denote the fibre class, and let $C_{0}$ be a section of minimal self-intersection. We write
$F^{2}=0,\qquad C_{0}\cdot F=1,\qquad C_{0}^{2}=-e,$
where $e\in\mathbb{Z}$ is the invariant of the ruled surface. For $B\cong\mathbb{P}^{1}$, one always has $e\geq 0$, and $X_{0}\cong\mathbb{F}_{e}$. For $g\geq 2$, one also has $e\geq 0$. When $g=1$, besides the cases $e\geq 0$, there is one exceptional indecomposable minimal ruled surface with e=1e=-1.
In all cases, the first Chern class is $c_{1}(X_{0})=2C_{0}+(2-2g+e)F$. A numerical class on $X_{0}$ can be written as $\omega]=aC_{0}+bF$, and one computes
@@
When $e\geq 0$, the Kähler cone is given by $a>0,\penalty 10000\ b>ae$. In the exceptional elliptic case $g=1$, e=1e=-1, the Kähler cone is given by $a>0,\penalty 10000\ b>-\frac{a}{2}$. Indeed, in this case the extra effective class $2C_{0}-F=-K_{X_{0}}$ appears, and positivity on $F$ and $2C_{0}-F$ is equivalent to the above condition.
@@
Let $X_{0}$ be a minimal ruled surface over a smooth complex curve $B$ of genus $g$, with invariant $e$. Then
@@
Moreover:
  1. (1)
    when $g=0$, the supremum is attained precisely on classes proportional to $C_{0}+(e+1)F$;
  2. (2)
    when $g=1$ and $e\geq 0$, the supremum is attained precisely when $b/a\geq e+1$;
  3. (3)
    when $g=1$ and e=1e=-1, the supremum is attained precisely when $b/a\geq 0$;
  4. (4)
    when $g>1$, the supremum is not attained.
@@
We split the proof according to the value of $e$.
Case 1: $e\geq 0$. In this case it is well known that the Mori cone $\overline{\operatorname{NE}}(X_{0})$ is generated by the numerical classes of $F$ and $C_{0}$. Hence
@@
Therefore
Since $\mathcal{J}_{X_{0}}$ is homogeneous of degree 0 in $\omega]$, it depends only on the ratio $t:=b/a$. The Kähler condition becomes $t>e$, and
$\mathcal{J}_{X_{0}}([\omega])=4\pi\,\phi_{e,g}(t),\qquad\phi_{e,g}(t):=\min\{1,\ t-e\}\cdot\frac{2t+2-2g-e}{2t-e},\qquad t>e.$
We now distinguish according to $g$.
Subcase 1a: $g\geq 1$. Since $2-2g\leq 0$, we have $2t+2-2g-e\leq 2t-e$, and hence
$\frac{2t+2-2g-e}{2t-e}\leq 1.$
Therefore
$\phi_{e,g}(t)\leq\min\{1,t-e\}\leq 1,$
so
$\sup_{t>e}\phi_{e,g}(t)\leq 1.$
If $g=1$, then $\phi_{e,1}(t)=\min\{1,t-e\}$, so $\sup_{t>e}\phi_{e,1}(t)=1$, and the supremum is attained precisely when $t-e\geq 1$, i.e. $b/a\geq e+1$. If $g>1$, then 22g<02-2g<0, so $\phi_{e,g}(t)<\min\{1,t-e\}\leq 1$ for all $t>e$. On the other hand,
$\lim_{t\to+\infty}\phi_{e,g}(t)=\lim_{t\to+\infty}\frac{2t+2-2g-e}{2t-e}=1.$
Hence $\sup_{t>e}\phi_{e,g}(t)=1$, but it is not attained. Thus, for all $g\geq 1$ and $e\geq 0$,
$\sup_{\mathcal{K}(X_{0})}\mathcal{J}_{X_{0}}([\omega])=4\pi.$
Subcase 1b: $g=0$. Now
For $t\geq e+1$,
$\phi_{e,0}^{\prime}(t)=\frac{d}{dt}\left(\frac{2t+2-e}{2t-e}\right)=-\frac{4}{(2t-e)^{2}}<0,$
so $\phi_{e,0}$ is strictly decreasing on $e+1,+\infty)$, and
$\max_{t\geq e+1}\phi_{e,0}(t)=\phi_{e,0}(e+1)=\frac{e+4}{e+2}.$
For $e<t\leq e+1$, a direct computation gives $\phi_{e,0}^{\prime}(t)>0$, so $\phi_{e,0}$ is strictly increasing on $(e,e+1]$, and
$\max_{e<t\leq e+1}\phi_{e,0}(t)=\phi_{e,0}(e+1)=\frac{e+4}{e+2}.$
@@
$\sup_{t>e}\phi_{e,0}(t)=\phi_{e,0}(e+1)=\frac{e+4}{e+2},$
and the supremum is attained precisely at $t=e+1$, i.e. on classes proportional to $C_{0}+(e+1)F$.
Case 2: $g=1$ and e=1e=-1. This is the exceptional indecomposable elliptic ruled surface. In this case the effective cone is generated by the classes $F,\penalty 10000\ C_{0},$ and $2C_{0}-F=-K_{X_{0}}$. Hence for a class $\omega]=aC_{0}+bF$, we have
$\omega]\cdot F=a,\qquad[\omega]\cdot C_{0}=a+b,\qquad[\omega]\cdot(2C_{0}-F)=a+2b.$
Then the Kähler condition becomes a>0,b>a/2a>0,\ b>-a/2. Writing $t=b/a$, with $t>-\frac{1}{2}$, we obtain
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])=a\min\{1,\ 1+2t\}.$
Moreover, $c_{1}(X_{0})=2C_{0}-F$, so
$c_{1}(X_{0})\cdot[\omega]=2b+a,\qquad[\omega]^{2}=a(2b+a).$
Therefore
$\mathcal{J}_{X_{0}}([\omega])=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\cdot 4\pi\frac{2b+a}{a(2b+a)}=4\pi\min\{1,\ 1+2t\}.$
It follows immediately that
$\sup_{\mathcal{K}(X_{0})}\mathcal{J}_{X_{0}}([\omega])=4\pi,$
and the supremum is attained precisely when $t\geq 0$ that is, $\frac{b}{a}\geq 0$. Combining the above cases proves the proposition. ∎
@@
When $B\cong\mathbb{P}^{1}$, the ruled surface $X_{0}$ is the Hirzebruch surface $\mathbb{F}_{e}$ with $e\geq 0$. In particular, if $e=1$, $\mathbb{F}_{1}$ is not minimal and is isomorphic to the blowup $\operatorname{Bl}_{1}\mathbb{P}^{2}$ of $\mathbb{P}^{2}$ at one point. As expected, our computations of $\mathcal{J}_{\operatorname{Bl}_{1}\mathbb{P}^{2}}$ in Example 3.5 and of $\mathcal{J}_{\mathbb{F}_{1}}$ in Proposition 4.2 yield the same value..
For later use, we also fix notation for blow-ups of ruled surfaces. Let $X_{k}=\operatorname{Bl}_{k}X_{0}$ be the blow-up of $X_{0}$ at $k$ points in very general position, and denote by $E_{1},\dots,E_{k}$ the exceptional divisors. Here “very general position” means that the blown-up points lie on distinct fibres of the ruling, and none of them lies on the distinguished section $C_{0}$. This is the worst configuration for our purposes: if some points lie on the same fibre, or on $C_{0}$, additional effective curves appear whose strict transforms can only decrease $\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$, and hence can only decrease $\mathcal{J}_{X_{k}}([\omega])$.
On $X_{k}$, $k\geq 1$, we have
@@
$C_{0}^{2}=-e,\qquad F^{2}=0,\qquad C_{0}\cdot F=1,$
and
$E_{i}^{2}=-1,\qquad E_{i}\cdot C_{0}=E_{i}\cdot F=E_{i}\cdot E_{j}=0\quad(i\neq j).$
We will write a Kähler class on $X_{k}$ in the form
$\omega]=aC_{0}+bF-\sum_{i=1}^{k}t_{i}E_{i},\qquad a>0,\quad t_{i}>0.$
The first Chern class is
@@
$\omega]^{2}=2ab-ea^{2}-\sum_{i=1}^{k}t_{i}^{2},$
and
$c_{1}(X_{k})\cdot[\omega]=2b+(2-2g-e)a-\sum_{i=1}^{k}t_{i}.$
These formulas will be used repeatedly in the subsequent subsections.
@@ 4.2. Non-rational ruled surfaces
Let $X_{0}$ be a minimal ruled surface over a smooth curve $B$ of genus $g\geq 1$, with invariant $e\geq-1$. For an integer $k\geq 1$, let $X_{k}=\text{Bl}_{k}(X_{0})$ be the blow-up at $k$ points in very general position, with exceptional divisors $E_{1},\dots,E_{k}$. Then for every $k\geq 1$ one has
@@
Since $\mathcal{J}_{X_{k}}([\omega])$ is homogeneous of degree 0 in $\omega]$, we may normalize $a=1$ and write
$\omega]=C_{0}+bF-\sum_{i=1}^{k}t_{i}E_{i},\qquad t_{i}>0,$
and $b>e$ if $e\geq 0$; b>1/2b>-1/2 if $g=1$ and e=1e=-1. In particular, since $\omega]$ is Kähler, we always have $\omega]^{2}=2b-e-\sum t_{i}^{2}>0$.
As in the minimal case, $\mathcal{J}_{X_{k}}([\omega])$ can be written as
@@
On the other hand, using $g\geq 1$ and $0<t_{i}<1$ we have
@@
It remains to show that the bound $2\pi$ is sharp. Fix $t_{1}=\cdots=t_{k}=1/2$ and consider the family of classes
$\omega_{b}]=C_{0}+bF-\frac{1}{2}\sum_{i=1}^{k}E_{i},\qquad b\gg 1.$
For $b$ sufficiently large, these classes lie in the Kähler cone.
We claim that, for all sufficiently large $b$,
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{b}])=\frac{1}{2}.$
Indeed, each exceptional divisor $E_{i}$ is effective and satisfies $\omega_{b}]\cdot E_{i}=1/2$. Moreover, since the blown-up points lie on distinct fibres, the strict transform of the fibre through $p_{i}$ is effective of class $F-E_{i}$, and $\omega_{b}]\cdot(F-E_{i})=1/2.$ Hence
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{b}])\leq\frac{1}{2}.$
Conversely, let $C\subset X_{k}$ be any irreducible effective curve distinct from the $E_{i}$ and the strict transforms $F-E_{i}$. Write its numerical class as
$C]=dC_{0}+nF-\sum_{i=1}^{k}m_{i}E_{i},\qquad d,n,m_{i}\in\mathbb{Z}_{\geq 0}.$
Then
$\omega_{b}]\cdot C=(b-e)d+n-\frac{1}{2}\sum_{i=1}^{k}m_{i}.$
If $d\geq 1$, this tends to $\infty$ as $b\to+\infty$, regardless of whether $e\geq 0$ or e=1e=-1. Thus it is strictly bigger than $1/2$ for all sufficiently large $b$. If $d=0$, then $C$ is supported in fibres and exceptional divisors. Since the blown-up points lie on distinct fibres, the only irreducible possibilities are $E_{i}$ and $F-E_{i}$, which have already been treated. Therefore every other irreducible effective curve has intersection strictly bigger than $1/2$ for $b\gg 1$, and hence
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{b}])=\frac{1}{2}.$
For these classes we have
$\mathcal{J}_{X_{k}}([\omega_{b}])=\frac{1}{2}\cdot 4\pi\,\frac{2b+2-2g-e-k/2}{2b-e-k/4}.$
As $b\to+\infty$, the scalar-curvature factor tends to $4\pi$, and therefore
$\mathcal{J}_{X_{k}}([\omega_{b}])\nearrow 2\pi.$
This proves
$\sup_{\mathcal{K}^{+}(X_{k})}\mathcal{J}_{X_{k}}([\omega])=2\pi.$
@@
Moreover, equality holds if and only if $B$ is an elliptic curve, $X$ is minimal and isometrically covered by $\mathbb{P}^{1}\times\mathbb{C}$, so that the induced metric $\omega$ has constant scalar curvature with holomorphic $2$-systole realized by the $\mathbb{P}^{1}$-fibre.
@@
By Theorem 2.2 (1), Proposition 4.2, and Lemma 4.4, every PSC Kähler metric $\omega$ on a ruled surface over a base of genus $g\geq 1$ satisfies
$\operatorname{sys}_{2}(\omega)\cdot\min_{X}S(\omega)\leq\mathcal{J}_{X}([\omega])\leq 4\pi.$
Moreover, if $X$ is non-minimal, then Lemma 4.4 gives the sharper bound
$\mathcal{J}_{X}([\omega])\leq 2\pi,$
so equality in
$\operatorname{sys}_{2}(\omega)\cdot\min_{X}S(\omega)\leq 4\pi$
can occur only when $X$ is minimal. By Proposition 4.2, the minimal case attains the value $4\pi$ only when $g=1$. Equality in Theorem 2.2 (1) then forces $\omega$ to have constant scalar curvature and the holomorphic $2$-systole to be realized by the $\mathbb{P}^{1}$-fibre.
It remains to prove that, in this equality case, $X$ is isometrically covered by $\mathbb{P}^{1}\times\mathbb{C}$. Since $X\xrightarrow{\pi}B$ is a minimal ruled surface over an elliptic curve, there exists a rank-$2$ holomorphic vector bundle $\mathcal{E}$ on $B$ such that $X=\mathbb{P}(\mathcal{E})$ and $\pi$ is the bundle projection (see [14, Chapter V]). By the classification of cscK metrics on ruled surfaces due to Apostolov and Tønnesen-Friedman [1, Theorem 2], the existence of a cscK metric on $X=\mathbb{P}(\mathcal{E})$ is equivalent to the slope-polystability of $\mathcal{E}$. Hence $\mathcal{E}$ is polystable. By Fujiki’s equivalences [10, Lemma 2], this implies that $X$ is quasi-stable. Equivalently, there exists a projective unitary representation
@@
where $\widetilde{B}\cong\mathbb{C}$ is the universal cover. Fix a biholomorphism $\Phi:X\stackrel{{\scriptstyle\sim}}{{\to}}X_{\rho}$, and set
@@
Then $\omega_{\rho}$ is a cscK Kähler metric on $X_{\rho}$, and $\Phi$ is an isometry between $(X,\omega)$ and $(X_{\rho},\omega_{\rho})$.
Since $g(B)=1$, Fujiki’s rigidity result [10, Lemma 10] shows that any cscK Kähler metric on $X_{\rho}$ is a generalized Kähler–Einstein metric in the sense of [10, §3]. It follows that $X_{\rho}$ is isometrically covered by $\mathbb{P}^{1}\times\mathbb{C}$ endowed with the product Kähler metric $\omega_{\mathrm{FS}}\oplus\omega_{\mathrm{flat}}$. Therefore $X$ is also isometrically covered by $\mathbb{P}^{1}\times\mathbb{C}$. ∎